# Starlight from JWST: Implications for star formation and dark matter models

### v2

@@ 3.2 Enhanced small-scale structures
For the enhancement of small-scale structures, we also consider two cases: 1) an axion DM model that includes axion miniclusters and 2) a scenario where heavy primordial black holes (PBHs) constitute a fraction of the DM density. In both cases, we model the enhancement of the matter power spectrum as (Hütsi et al. 2023)
@@
This parametrization includes one parameter, the scale $k_{c}$ above which the power spectrum is dominated by a white noise contribution. We perform the analysis in terms of $k_{c}$, which is related to the parameters of the underlying DM models as discussed below.
In axion-like particle DM models, where the global $U(1)$ symmetry is broken after cosmic inflation, significant small-scale density fluctuations arise due to the Kibble mechanism. These lead to the formation of axion miniclusters (Kolb & Tkachev 1993). This scenario results in a white noise contribution to the matter power spectrum (Fairbairn et al. 2018; Feix et al. 2019; Ellis et al. 2022). The resulting power spectrum can be approximated by Eq. (14) with
@@
where $m_{a}$ denotes the axion-like particle mass. The white noise contribution is cut at scales that re-entered horizon when the axion began oscillating:
@@
In scenarios involving heavy PBHs there is a white noise contribution to the matter power spectrum (Inman & Ali-Haïmoud 2019; De Luca et al. 2020) with
@@
where $f_{\rm PBH}$ is the DM fraction in PBHs, $m_{\rm PBH}$ is the PBH mass, and $\rho_{\rm DM}$ is the dark matter density. The cut-off is set by the average separation of PBHs (Hütsi et al. 2023):
$k_{\rm cut}=900h\,\mathrm{Mpc}^{-1}\left[f_{\rm PBH}\frac{10^{4}M_{\odot}}{m_{\rm PBH}}\right]^{1/3}\,.$ (18)
Below this scale, one expects roughly one PBH per corresponding comoving volume, and the seed effect (Carr & Silk 1983, 2018; Cappelluti et al. 2022) becomes dominant. The condition $k_{c}<k_{\rm cut}$ is satisfied for $f_{\rm PBH}>10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}$.
@@ 3.3 Window function
$\sigma^{2}(R)=A\int{\rm d}k\,k^{2}P(k)\tilde{W}^{2}(k|R)\,,$ (19)
where the normalization factor $A$ is chosen to match the Planck measurement of $\sigma_{8}=\sigma(R=8/h\,{\rm Mpc})=0.811$ (Aghanim et al. 2020). We use a window function of the form (Leo et al. 2018; Verwohlt et al. 2024):
@@
where $R$ is related to the mass of the overdensity by $M=4\pi\bar{\rho}R^{3}/3$. In the case of a suppressed matter power spectrum, the HMF scales as ${\rm d}n/{\rm d}\ln M\propto M^{c_{2}/6}$ and the halo growth rate as ${\rm d}M/{\rm d}z\propto M^{1-c_{2}/6}$ at scales much below the suppression scale. For all models, we use $c_{2}=6$ so that also for WDM and FDM models the growth rate increases with the halo mass, and fix $c_{1}=0.43$ so that the resulting variance in CDM matches well with that computed using the standard real space top hat window function, in agreement with simulation results. We find that this choice also reproduces well the simulation results in the FDM model (Schive et al. 2016), but it underestimates the halo mass where the HMF found in WDM simulations deviates from the CDM prediction (Lovell et al. 2014). On the other hand, the choice $c_{1}=1/3.3$ suggested by Leo et al. (2018) reproduces the low-mass part, but it does not match the CDM prediction at high masses. To reproduce the simulation results in WDM and to match CDM predictions at masses much above the suppression scale, we use $c_{1}=0.43$ in the window function and multiply $\alpha$ in the WDM transfer function by $3.3c_{1}\approx 1.4$.
@@
The HMFs are shown in Fig. 3, where they are compared with the CDM predictions. For the case with suppressed small-scale structures, the HMF grows $\propto M$ at small masses. This power-law behaviour does not affect significantly the fits to the UV luminosity observations discussed below. For the enhanced perturbations, we can see how the enhancement gets weaker at low $z$ as the miniclusters merge with heavier halos and their original imprint is diluted.
@@ 4 UV luminosity functions
where ${\rm d}P(L|M)/{\rm d}L$ denotes the distribution of luminosities emitted by galaxies in halos of mass $M$ and ${\rm d}P(L_{\rm obs}|L)/{\rm d}L_{\rm obs}$ is the distribution of observed luminosities if the emitted luminosity is $L$. The latter is affected by gravitational lensing magnification (Takahashi et al. 2011) and dust attenuation (Wang & Heckman 1996):
@@ 4.1 Lensing magnification
The first correction to the observed luminosity that we consider is gravitational lensing. This correction is relevant for the bright end of the UV luminosity function, which receives contributions from magnified distant galaxies lensed by intervening structures. As shown by Takahashi et al. (2011), the magnification can be approximated by [^4] This approximation underestimates the high-$\mu$ tail of the distribution. However, this tail is damped by the finite sizes of the lenses.
@@
We compute the distribution of $\mu$ by generating numerically the distribution of $\kappa$ and converting it to ${\rm d}P(\mu)/{\rm d}\mu$ using Eq. (23). We model the lenses by NFW halos whose spatial distribution follows a uniform distribution and use the halo mass-concentration/redshift relation derived by Ludlow et al. (2016). We account for the finite source sizes by averaging the surface density over a sphere of radius $R_{s}$ that we obtain by projecting the source size to the lens plane. We fix the source size to $10\,{\rm kpc}$ (see in Takahashi et al. (2011); Ferrami & Wyithe (2023) how the result is affected by the source size). The distribution of lenses is given by
@@
We compute the distribution of the convergence by picking random numbers from ${\rm d}P^{(1)}(\kappa)/{\rm d}\kappa$ to generate multiple realizations of $\sum_{j}\kappa^{(1)}(\boldsymbol{\theta}_{j})$ and shifting the resulting distribution by $\kappa_{\rm E}$ so that its mean is zero. [^5] A similar approach of finding the distribution of a sum of random variables was introduced by Ellis et al. (2024) for the computation of the gravitational wave background from a population of supermassive black hole binaries.
@@
Figure 5: The UV luminosity function at $z=4$ (black) and at $z=17$ (blue) with and without corrections due to lensing magnification and dust attenuation. The cyan and pink points show, respectively, the HST $z=4$ data from Bouwens et al. (2021) and Harikane et al. (2022), and the brown and gray points the JWST $z=17$ data from Pérez-González et al. (2025) and Castellano et al. (2025).
@@ 4.2 Dust attenuation
Following the literature (see, e.g., Trenti et al. (2015); Vogelsberger et al. (2020)), we neglect the variance in the dust attenuation $B$ and approximate ${\rm d}P(B)/{\rm d}B\approx\delta(B-\bar{B})$. This gives
@@
The UV luminosity functions are commonly presented in terms of the absolute magnitude instead of the luminosity. The associated absolute magnitude $M_{\rm UV}$ is defined by (Oke & Gunn 1983)
@@
The latter is obtained by fitting the results shown in Table 3 of Bouwens et al. (2014), assuming that $\bar{\beta}_{\rm UV}$ scales exponentially with $M_{\rm UV}$. Notice that, in Eq. (29), instead a piecewise function, we use a softplus function with sharpness parameter $s=3$ chosen so that the resulting dust attenuation roughly matches that obtained by numerical sampling shown in Fig. 1 of Vogelsberger et al. (2020).
@@
where $M_{\rm UV}=M_{\rm obs}-A_{\rm UV}+1.086\ln\mu$. Comparing the short- and long-dashed curves in Fig. 5, we see that dust attenuation suppresses strongly the high-luminosity tail of the luminosity function at low redshifts but at very high redshifts its effect is almost negligible. However, we note that the dust attenuation is calibrated to observations at $z\leq 8$ (Bouwens et al. 2014) and we have simply extrapolated these results to $z\gg 8$.
@@ 4.3 Star formation rate
As mentioned in the Introduction, we assume that the SFR is proportional to the halo growth rate (Bian et al. 2013), $\dot{M}_{*}=f_{B}f_{*}(M)\,\dot{M}$, where $f_{B}=\Omega_{B}/\Omega_{M}\approx 0.16$. We parametrize the proportionality coefficient $f_{*}(M)$, which models the feedback effects from SNe and AGNs, as a broken power-law around $M=M_{c}$ with an exponential suppression in halos with $M<M_{t}<M_{c}$:
@@
where $\alpha,\,\beta,\,\epsilon>0$. A similar parametrization is used, e.g., in the GALLUMI code (Sabti et al. 2022a). The difference is that, in our parametrization, the maximum of $f_{*}(M)$ is at $M=M_{c}$ for $M_{t}\ll M_{c}$. We find that $z$-independent values of $\alpha$, $\beta$, $\epsilon$ and $\kappa_{\rm UV}$ give a good fit to the observations up to $z\approx 10$, but that the data prefer higher luminosities at $z\gtrsim 10$.
To accommodate this enhancement in the luminosities of the early stars, we introduce a parametrization in which the conversion factor $\kappa_{\rm UV}$ changes around $z=z_{\kappa}$:
@@
where $\kappa_{0}=1.15\times 10^{-22}M_{\odot}\,{\rm s}{\rm erg}^{-1}\,{\rm yr}^{-1}$, $0<f_{\kappa}\leq 1$ and $\gamma>0$. The conversion factor depends on the initial stellar mass function and the value of $\kappa_{0}$ corresponds to the Salpeter initial mass function (Madau & Dickinson 2014). We allow for a wide range of possibilities in the transition to Pop-III stars; $z_{\kappa}$ determines the redshift at which the change of $\kappa_{\rm UV}$ starts, $\gamma$ parametrises how sharp the transition is and $f_{\kappa}$ the fractional change in $\kappa$, such that at $z\gg z_{\kappa}$, $\kappa_{\rm UV}=f_{\kappa}\kappa_{0}$. We show in the upper left panel of Fig. 6 $\kappa_{\rm UV}$ as a function of $z$ for the best fit values in the CDM model.
In order to fit the highest-$z$ UV data, we find that the feedback effects, which we parametrise with $f_{*}(M)$, need to be reduced. To allow for such a reduction, we parametrize the powers $X=\alpha,\beta$ so that they can approach zero linearly with $z$ above some redshift $z_{\rm fb}$:
@@
In this parametrization, $\alpha$ and $\beta$ become zero at $z>z_{*}$, and $f_{*}(M)$ is mass independent at $M\gg M_{t}$, as shown in the upper right panel of Fig. 6 for the best fit values. This implies that in all halos with $M\gg M_{t}$, the baryons that the halo accretes are converted into stars with efficiency $\epsilon$.
@@
Figure 7: The UV luminosity function at different redshifts in the cold (black), fuzzy (green) and warm (orange) DM models and in the model with enhanced matter power (red). The best fit is shown for CDM and WDM cases, while for FDM the mass corresponds to the 95% lower bound. The dashed and dot-dashed curves show the CDM UV luminosity function without the suppression of the feedback effects and without the change in the UV conversion factor $\kappa_{\rm UV}$. The points with errorbars show the HST measurements from Bouwens et al. (2021) (cyan) and Harikane et al. (2022) (pink) and the JWST measurements from Donnan et al. (2024) (blue), Pérez-González et al. (2025) (brown) and Castellano et al. (2025) (gray).
@@ 5 Results
We consider UV luminosity function measurements derived from both HST observations (Bouwens et al. 2021; Harikane et al. 2022) and JWST observations (Donnan et al. 2024; Pérez-González et al. 2025; Castellano et al. 2025), and perform a Markov-chain Monte Carlo (MCMC) analysis of the models. For each scan over the model parameters, we generate 10 MCMC chains of 40000 steps, including 10000 burn-in steps, using the Metropolis-Hastings sampler with Gaussian proposal distributions whose widths we choose so that the acceptance rate is around $10\%$. The model parameter priors are given in Table 1. We assume that the data follow a split normal distribution $\mathcal{N}(\mu,\sigma_{1},\sigma_{2})$, where $\mu$ denotes the mode and $\sigma_{1,2}$ the left- and right-hand-side standard deviations. The likelihood is given by
@@
$M_{t}/M_{\odot}$ LogU$(6,10)$ $10^{7.4}$
$M_{c}/10^{11}M_{\odot}$ U$(3,5)$ 3.9
$\epsilon$ U$(0.37,0.43)$ 0.39
$\alpha$ U$(0.6,1.1)$ 0.88
$\beta$ U$(0.2,0.6)$ 0.40
$\gamma$ U$(0.05,0.6)$ 0.21
$f_{\kappa}$ U$(0.05,0.7)$ 0.43
$z_{\kappa}$ U$(9.8,11.4)$ 10.6
$z_{\rm fb}$ U$(7,25)$ 10.5
$z_{*}$ U$(20,36)$ 26.2
$\sigma_{\rm UV}$ U$(0.05,0.2)$ 0.055
Table 1: Prior ranges and best-fit values of the parameters in the CDM model.

### v4

@@ 3.2 Enhanced small-scale structures
For the enhancement of small-scale structures, we also consider two cases: 1) an axion DM model that includes axion miniclusters and 2) a scenario where heavy primordial black holes (PBHs) constitute a fraction of the DM density. In both cases, we model the enhancement of the matter power spectrum as (Hutsi:2022fzw)
@@
This parametrisation includes one parameter, the scale $k_{c}$ above which the power spectrum is dominated by a white noise contribution. We perform the analysis in terms of $k_{c}$, which is related to the parameters of the underlying DM models as discussed below.
In axion-like particle DM models, where the global $U(1)$ symmetry is broken after cosmic inflation, significantly small-scale density fluctuations arise due to the Kibble mechanism. These lead to the formation of axion miniclusters (Kolb:1993zz). This scenario results in a white noise contribution to the matter power spectrum (Fairbairn:2017sil; Feix:2019lpo; Ellis:2022grh). The resulting power spectrum can be approximated by Eq. (14) with
@@
where $m_{a}$ denotes the axion-like particle mass. The white noise contribution is cut at scales that re-entered the horizon when the axion began oscillating:
@@
In scenarios involving heavy PBHs there is a white noise contribution to the matter power spectrum (Inman:2019wvr; DeLuca:2020jug) with
@@
where $f_{\rm PBH}$ is the DM fraction in PBHs, $m_{\rm PBH}$ is the PBH mass, and $\rho_{\rm DM}$ is the dark matter density. The cut-off is set by the average separation of PBHs (Hutsi:2022fzw):
[…]
Below this scale, one expects roughly one PBH per corresponding comoving volume, and the seed effect (1983ApJ...268....1C; Carr:2018rid; Cappelluti:2021usg) becomes dominant. The condition $k_{c}<k_{\rm cut}$ is satisfied for $f_{\rm PBH}>10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}$.
@@ 3.3 Window function
$\sigma^{2}(R)=A\int{\rm d}k\,k^{2}\mathcal{P}(k)\tilde{W}^{2}(k|R)\,$, (19)
where the normalization factor $A$ is chosen to match the Planck measurement of $\sigma_{8}=\sigma(R=8/h\,{\rm Mpc})=0.811$ (Planck:2018vyg). We use a window function of the form (Leo:2018odn; Verwohlt:2024efh):
@@
where $R$ is related to the mass of the overdensity by $M=4\pi\bar{\rho}R^{3}/3$. In the case of a suppressed matter power spectrum, the HMF scales as ${\rm d}n/{\rm d}\ln M\propto M^{c_{2}/6}$ and the halo growth rate as ${\rm d}M/{\rm d}z\propto M^{1-c_{2}/6}$ at scales much below the suppression scale. For all models, we use $c_{2}=6$ so that also for WDM and FDM models the growth rate increases with the halo mass, and fix $c_{1}=0.43$ so that the resulting variance in CDM matches well with that computed using the standard real space top hat window function, in agreement with simulation results. We find that this choice also reproduces well the simulation results in the FDM model (Schive:2015kza), but it underestimates the halo mass where the HMF found in WDM simulations deviates from the CDM prediction (Lovell:2013ola). On the other hand, the choice $c_{1}=1/3.3$ suggested by Leo:2018odn reproduces the low-mass part, but it does not match the CDM prediction at high masses. To reproduce the simulation results in WDM and to match CDM predictions at masses much above the suppression scale, we use $c_{1}=0.43$ in the window function and multiply $\alpha$ in the WDM transfer function by $3.3c_{1}\approx 1.4$.
@@
The HMFs are shown in Fig. 3, where they are compared with the CDM predictions. For the case with suppressed small-scale structures, the HMF grows $\propto M$ at small masses. This power-law behaviour does not significantly affect the fits to the UV luminosity observations discussed below. For the enhanced perturbations, we can see how the enhancement gets weaker at low $z$ as the miniclusters merge with heavier halos and their original imprint is diluted.
@@ 4 UV luminosity functions
where ${\rm d}P(L|M)/{\rm d}L$ denotes the distribution of luminosities emitted by galaxies in halos of mass $M$ and ${\rm d}P(L_{\rm obs}|L)/{\rm d}L_{\rm obs}$ is the distribution of observed luminosities if the emitted luminosity is $L$. The latter is affected by gravitational lensing magnification (Takahashi:2011qd) and dust attenuation (1996ApJ...457..645W):
@@ 4.1 Lensing magnification
The first correction to the observed luminosity that we consider is gravitational lensing. This correction is relevant for the bright end of the UV luminosity function, which receives contributions from magnified distant galaxies lensed by intervening structures. As shown by Takahashi:2011qd, the magnification can be approximated by [^4] This approximation underestimates the high-$\mu$ tail of the distribution. However, this tail is damped by the finite sizes of the lenses.
@@
We compute the distribution of $\mu$ by generating numerically the distribution of $\kappa$, converting it to ${\rm d}P(\mu)/{\rm d}\mu$ using Eq. (23) and applying a $1/\mu$ factor to transform the image-plane distribution into the source-plane distribution. We model the lenses by NFW halos whose spatial distribution follows a uniform distribution and use the halo mass-concentration/redshift relation derived by Ludlow:2016ifl. We account for the finite source sizes by averaging the surface density over a sphere of radius $R_{s}$ that we obtain by projecting the source size to the lens plane. We fix the source size to $10\,{\rm kpc}$ (see in Takahashi:2011qd; Ferrami:2022xfy how the result is affected by the source size). The distribution of lenses is given by
@@
We compute the distribution of the convergence by picking random numbers from ${\rm d}P^{(1)}(\kappa)/{\rm d}\kappa$ to generate multiple realizations of $\sum_{j}\kappa^{(1)}(\boldsymbol{\theta}_{j})$ and shifting the resulting distribution by $\kappa_{\rm E}$ so that its mean is zero. [^5] A similar approach of finding the distribution of a sum of random variables was introduced by Ellis:2023dgf for the computation of the gravitational wave background from a population of supermassive black hole binaries.
@@
Figure 5: The UV luminosity function at $z=4$ (black) and at $z=17$ (blue) with and without corrections due to lensing magnification and dust attenuation. The cyan and pink points show, respectively, the HST $z=4$ data from 2021AJ....162...47B and 2022ApJS..259...20H, and the brown and gray points the JWST $z=17$ data from Perez-Gonzalez:2025bqr and Castellano:2025vkt.
@@ 4.2 Dust attenuation
Following the literature (see, e.g., Trenti:2014hka; 2020MNRAS.492.5167V), we neglect the variance in the dust attenuation $B$ and approximate ${\rm d}P(B)/{\rm d}B\approx\delta(B-\bar{B})$. This gives
@@
The UV luminosity functions are commonly presented in terms of the absolute magnitude instead of the luminosity. The associated absolute magnitude $M_{\rm UV}$ is defined by (Oke:1983nt)
@@
The latter is obtained by fitting the results shown in Table 3 of Bouwens:2013hxa, assuming that $\bar{\beta}_{\rm UV}$ scales exponentially with $M_{\rm UV}$. Notice that, in Eq. (29), instead a piecewise function, we use a softplus function with sharpness parameter $s=3$ chosen so that the resulting dust attenuation roughly matches that obtained by numerical sampling shown in Fig. 1 of 2020MNRAS.492.5167V.
@@
where $M_{\rm UV}=M_{\rm obs}-A_{\rm UV}+1.086\ln\mu$. Comparing the short- and long-dashed curves in Fig. 5, we see that dust attenuation suppresses strongly the high-luminosity tail of the luminosity function at low redshifts but at very high redshifts its effect is almost negligible. However, we note that the dust attenuation is calibrated to observations at $z\leq 8$ (Bouwens:2013hxa) and we have simply extrapolated these results to $z\gg 8$.
@@ 4.3 Star formation rate
As mentioned in the Introduction, we assume that the SFR is proportional to the halo growth rate (2013ApJ...774...28B), $\dot{M}_{*}=f_{B}f_{*}(M)\,\dot{M}$, where $f_{B}=\Omega_{B}/\Omega_{M}\approx 0.16$. We parametrise the proportionality coefficient $f_{*}(M)$, which models the feedback effects from SNe and AGNs, as a broken power-law around $M=M_{c}$ with an exponential suppression in halos with $M<M_{t}<M_{c}$:
@@
where $\alpha,\,\beta,\,\epsilon>0$. A similar parametrization is used, e.g., in the GALLUMI code (Sabti:2021xvh). The difference is that, in our parametrisation, the maximum of $f_{*}(M)$ is at $M=M_{c}$ for $M_{t}\ll M_{c}$. We find that $z$-independent values of $\alpha$, $\beta$, $\epsilon$ and $\kappa_{\rm UV}$ give a good fit to the observations up to $z\approx 10$, but that the data prefer higher luminosities at $z\gtrsim 10$.
To accommodate this enhancement in the luminosities of the early stars, we introduce a parametrisation in which the conversion factor $\kappa_{\rm UV}$ changes around $z=z_{\kappa}$:
@@
where $\kappa_{0}=1.15\times 10^{-22}M_{\odot}\,{\rm s}\,{\rm erg}^{-1}\,{\rm Myr}^{-1}$, $0<f_{\kappa}\leq 1$ and $\gamma>0$. The conversion factor depends on the initial stellar mass function and the value of $\kappa_{0}$ corresponds to the Salpeter initial mass function (Madau:2014bja). We allow for a wide range of possibilities in the transition to Pop-III stars; $z_{\kappa}$ determines the redshift at which the change of $\kappa_{\rm UV}$ starts, $\gamma$ parametrises how sharp the transition is and $f_{\kappa}$ the fractional change in $\kappa$, such that at $z\gg z_{\kappa}$, $\kappa_{\rm UV}=f_{\kappa}\kappa_{0}$. We show in the upper left panel of Fig. 6 $\kappa_{\rm UV}$ as a function of $z$ for the best fit values in the CDM model.
In order to fit the highest-$z$ UV data, we find that the feedback effects, which we parametrise with $f_{*}(M)$, need to be reduced. To allow for such a reduction, we parametrise the powers $X=\alpha,\beta$ so that they can approach zero linearly with $z$ above some redshift $z_{\rm fb}$:
@@
In this parametrisation, $\alpha$ and $\beta$ become zero at $z>z_{*}$, and $f_{*}(M)$ is mass independent at $M\gg M_{t}$, as shown in the upper right panel of Fig. 6 for the best fit values. This implies that in all halos with $M\gg M_{t}$, the baryons that the halo accretes are converted into stars with efficiency $\epsilon$.
@@ 5 Results
We consider UV luminosity function measurements derived from both HST observations (2021AJ....162...47B; 2022ApJS..259...20H) and JWST observations (2024MNRAS.533.3222D; Perez-Gonzalez:2025bqr; Castellano:2025vkt). The latter are based on high-$z$ photometric measurements, most of which are not spectroscopically confirmed. This can lead to errors in the estimation of redshift. Most notably, a photometric measurement of CEERS-93316 indicated that this galaxy was at $z\approx 16$ (2022arXiv220802794N) but a later spectroscopic measurement revealed it to be at $z\approx 5$ (2023Natur.622..707A). However, the redshifts of some of the high-$z$ galaxies have been confirmed spectroscopically by 2024MNRAS.533.3222D.
We perform a Markov-chain Monte Carlo (MCMC) analysis of the models. For each scan over the model parameters, we generate 8 MCMC chains, each consisting of 12000 samples of which the first 2000 are discarded as burn-in. We use the Metropolis-Hastings sampler with Gaussian proposal distributions whose widths we choose so that the acceptance rate is around $10\%$. The model parameter priors are given in Table 6. We assume that the data follow a split normal distribution $\mathcal{N}(\mu,\sigma_{1},\sigma_{2})$, where $\mu$ denotes the mode and $\sigma_{1,2}$ the left- and right-hand-side standard deviations. The likelihood is given by
@@
Table 1: Parameters range and best fit values [^6] Prior ranges and best-fit values of the parameters in the CDM model. The parametrization is given by Eqs. (32), (34), (35) and (36).
@@
$M_{t}/M_{\odot}$ LogU(6,10)(6,10) $10^{7.9}$
$M_{c}/10^{11}M_{\odot}$ U(3,5)(3,5) 4.0
$\epsilon$ U(0.36,0.42)(0.36,0.42) 0.39
$\alpha$ U(0.6,1.1)(0.6,1.1) 0.88
$\beta$ U(0.2,0.6)(0.2,0.6) 0.40
$\gamma$ U(0.05,0.6)(0.05,0.6) 0.27
$f_{\kappa}$ U(0.05,0.7)(0.05,0.7) 0.29
$z_{\kappa}$ U(9.8,11.4)(9.8,11.4) 10.7
$z_{\rm fb}$ U(7,25)(7,25) 12.7
$z_{*}$ U(20,36)(20,36) 22.9
$\sigma_{\rm UV}$ U(0.05,0.2)(0.05,0.2) 0.068
Figure 7: Posteriors of SFR fits to the UV luminosity function observations in different models. The contours indicate the 68% and 95% credible regions. The marginalised 68% credible ranges and the 95% credible constraints are shown for CDM on top of each column.